# Reflection of Nichols Algebras over Coquasi-Hopf Algebras

### v1

@@ 2.4 Nichols algebras in braided monoidal category
Let $V$ be a nonzero object in $\mathcal{C}$, by $T(V)$ we denote the tensor algebra in $\mathcal{C}$ generated freely by $V$. Here we denote that
@@
[defofnichols, Definition 2.4] For a given object $V\in\mathcal{C}$ the Nichols algebra $\mathcal{B}(V)$ is the unique Hopf algebra in $\mathcal{C}$ that satisfies the following conditions:
@@
Recall that the Nichols algebra in $\mathcal{C}$ are of the form $T(X)/I$ where $X\in\mathcal{C}$ and $I$ is the unique maximal graded Hopf ideal in $T(V)$ generated by homogeneous elements of degree greater than or equal to $2$. According to $F(T(V))\cong T(F(V))$, we have
@@
which is finite-dimensional as well. Here $F(I)$ is a homogeneous Hopf ideal of degree greater than or equal to $2$ of $T(F(V))$. Note that Nichols algebra generated by $F(V)$ must be of the form $T(F(V))/J$, where $J$ is the unique maximal homogeneous graded Hopf ideal of $T(F(V))\in\mathcal{D}$ with degree greater than or equal to $2$. Hence $F(I)\subset J$ and $T(F(V))/J\subset T(F(V))/F(I)$. On the other hand, $F^{-1}:\mathcal{D}\longrightarrow\mathcal{C}$ is inverse of $F$, which is an exact monoidal functor. Hence
@@ 3 Monoidal equivalence between Yetter-Drinfeld module categories
We build upon the theory of Hopf pairings and partial dualization in braided monoidal categories Simon. We apply this theory to the setting of coquasi-Hopf algebras. Our main result, Theorem 3.8, establishes a key braided monoidal equivalence $\Omega$ between categories of Yetter-Drinfeld modules, which is important to our reflection theory.
@@ 3.1 Hopf pairings in braided monoidal categories
In this subsection, $\mathcal{C}$ always represents a braided monoidal category with associative isomorphism $a$ and braiding isomorphism $c$ except for special instructions. We begin by introducing the notion of Hopf pairings.
@@
A Hopf pairing $\omega:A\otimes B\to 1$ is called non-degenerate, if there is a morphism $\omega^{\prime}:1\rightarrow B\otimes A$, such that
@@
[Simon, Theorem 3.20] Let $\omega:A\otimes B\to\mathbf{1}$ be a non-degenerate Hopf pairing. The categories $^{A}_{A}\mathcal{YD}(\mathcal{C})$} and $^{B}_{B}\mathcal{YD}(\mathcal{C})$} are braided monoidal equivalent via the monoidal functors $\Omega^{\omega}$ as follows.
@@
    is isomorphic to the identity functor by [Simon, Proposition 3.22].
@@
    For completeness, we recall the construction of the functor $\Omega^{\omega}$, which is investigated in [Simon, Theorem 3.16, Corollary 3.19]. Let $\omega:A\otimes B\to\mathbf{1}$ be a non-degenerate Hopf pairing. The functor $\Omega^{\omega}$ is given by
@@ 3.2 Partial dualization of a Hopf algebra in braided monoidal categories
We now turn our attention to the construction of partial dualization. This construction, originally developed in [Simon, Theorem 4.4], provides a powerful tool for studying Hopf algebras in braided monoidal categories.
@@
We establish this result in several steps. First, the object $\mathcal{B}(N)^{*\operatorname{op,cop}}$ with the following operations is a Hopf algebra in $^{H}_{H}\mathcal{YD}$} by [AGn99, Lemma 2.2.3]
@@
in $^{H}_{H}\mathcal{YD}$}. The object $\mathcal{B}(N^{*})$ is the graded dual of $\mathcal{B}(N)$ by [defofnichols, Lemma 2.6]. Therefore $\operatorname{dim}(\mathcal{B}(N^{*}))=\operatorname{dim}(\mathcal{B}(N)^{*})$ since $\operatorname{dim}(\mathcal{B}(N))$ is finite-dimensional. As a result
@@
Let $H$ be a coquasi-Hopf algebra with bijective antipode, and $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. Suppose $N=\bigoplus_{i\in I}N_{i}\in{{}^{H}_{H}\mathcal{YD}}$, and $\mathcal{B}(N_{i})$ are all finite-dimensional. For each $i\in I$, we have a partial dualization datum $\mathcal{A}_{i}=(\mathcal{B}(N)\overset{\pi_{i}}{\rightarrow}\mathcal{B}(N_{i}),\mathcal{B}(N_{i}^{*}),\operatorname{ev}_{\mathcal{B}(N_{i})}^{-})$. We set $K_{i}:=\mathcal{B}(N)^{\operatorname{co}\mathcal{B}(N_{i})}\in{{}^{\mathcal{B}(N_{i})}_{\mathcal{B}(N_{i})}\mathcal{YD}(\mathcal{C})}$. Then we have the following equivalences of braided monoidal categories
@@ 3.3 The functor $\Omega^{\omega}$ under $\mathbb{Z}$-gradings
(1) The module and comodule maps of $K(1)$ are $\mathbb{Z}$-graded and hence the infinitesimal braiding $c\in\mathrm{Aut}(K(1)\otimes K(1))$, being determined by the module and comodule maps, is $\mathbb{Z}$-graded. Moreover, the structure map of $K(n)$ for $n\in\mathbb{N}$ are determined by $c$ and $K(1)$, since $K(1)$ generates $K$. Therefore $K$ is a $\mathbb{Z}$-graded objects in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$.
@@
Let $H$ be a coquasi-Hopf algebra with bijective antipode, and $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. Suppose $R$ is an $\mathbb{N}_{0}$-graded Hopf algebra in $^{H}_{H}\mathcal{YD}$}. Let $X$ be a $R$-module in $\mathcal{C}$ and $Y$ be a $R$-comodule in $\mathcal{C}$. For $n\geq 0$, we define:
@@
With above notations, the functor $\Omega$ relates the operator $\mathcal{F}_{n}$ to $\mathcal{F}^{n}$.
@@
Define an auxiliary map $\rho^{\prime}:=(\operatorname{ev}\otimes\operatorname{id}_{V})\circ a^{-1}_{\mathcal{B}(N^{*}),\mathcal{B}(N),V}\circ(\operatorname{id}_{\mathcal{B}(N^{*})}\otimes\delta_{\mathcal{B}(N)})$, which is derived from $\rho_{\mathcal{B}(N^{*})}$. By [rootsys, Corollary 3.3.4, Lemma 12.2.11], $(V,\rho^{\prime})$ is still a $\mathcal{B}(N^{*})$-module in $\mathcal{C}$ and
@@
is $\bigoplus_{m\neq n}\mathcal{B}(N^{*})(m)\otimes V$. This means, if $v\in\mathcal{F}^{n}V$, then for each $i>n$, $\rho^{\prime}(\mathcal{B}(N^{*})(i)\otimes v)=0$ by (3.5). On the other hand, if $v\in\mathcal{F}_{n}(V,\rho^{\prime})$, this will imply $\delta(v)\in\bigoplus_{i=0}^{n}\mathcal{B}(N)(i)\otimes V$. which completes the proof of (1): $\mathcal{F}_{n}\Omega(V)=\mathcal{F}_{n}(V,\rho^{\prime})=\mathcal{F}^{n}V$.
@@
We denote $\delta^{\prime}=(\operatorname{id}_{\mathcal{B}(N^{*})}\otimes\rho_{\mathcal{B}(N)})\circ a_{{\mathcal{B}(N^{*})},{\mathcal{B}(N)},V}\circ((\operatorname{ev}^{-})^{\prime}\otimes\operatorname{id}_{V})$, which origins from $\delta_{\mathcal{B}(N)}$. Again by [rootsys, Corollary 3.3.6, Lemma 12.2.11]
@@ 4 Projections of Nichols algebras
Let $(\mathbbm{k}G,\Phi)$ be a pointed coquasi-Hopf algebra, where $G$ is a finite group, $\Phi$ is a $3$-cocycle on $G$. In this section, we study projection of Nichols algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}, which will play an important role when considering reflection of simple Yetter-Drinfeld modules. The proofs of the some lemmas and theorems in this section, while are conceptually parallel to those in the Hopf algebra setting HSdualpair, takes a different path due to the non-strict monoidal structure.
@@ 4.1 The structure of the space of coinvariant $K$
To study the projection properties of Nichols algebras, we must first examine their structure in the cateogry $^{G}_{G}\mathcal{YD}^{\Phi}$}.
@@ 4.2 The semisimplicity of the object $L$
[rootsys, Lemma 1.3.13, Proposition 1.3.14]
@@
[rootsys,Proposition 13.2.3 Let $C$ be an $\mathbb{N}_{0}$-graded coalgebra, $X$ an $\mathbb{N}_{0}$-graded left $C$-comodule, and $Y$ a $C$-subcomodule of $X$. Let $k>0$ be an integer. Assume that the map $\delta_{n-k,k}^{X}:X(n)\to C(n-k)\otimes X(k)$ is injective for all $n\geq k$, and $Y$ is not contained in $\bigoplus_{i=0}^{k-1}X(i)$. Then $Y\cap\bigoplus_{i=0}^{k}X(i)\neq 0$.
@@
Now let $X\in L$, then $\delta(X)\in\mathcal{B}(N)\otimes L$ by Lemma 4.4, which implies that L inherits an $\mathbb{N}_{0}$-graded $\mathcal{B}(N)$-comodule via the restriction of $\delta:L\rightarrow\mathcal{B}(N)\otimes L$.
@@ 4.3 Realizing $K$ as a Nichols algebra
We endow a non-standard grading on $\mathcal{B}(M\oplus N)$ via setting $\operatorname{deg}(M)=1$ and $\operatorname{deg}(N)=0$. Under this grading, $\mathcal{B}(M\oplus N)$ remains a $\mathbb{N}_{0}$-graded Hopf algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}. Furthermore, $\mathcal{B}(M\oplus N)\#\mathbbm{k}G$ becomes a $\mathbb{N}_{0}$-graded coquasi-Hopf algebra with $\operatorname{deg}(\mathbbm{k}G)=0$. It follows direct that $K$ inherits an $\mathbb{N}_{0}$-graded Hopf algebra in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with $K(n)=K\cap(\mathcal{B}(M\oplus N)\#\mathbbm{k}G)(n)$. In particular, we have $K(1)=K\cap(\mathcal{B}(M\oplus N)\#\mathbbm{k}G)(1)=K\cap(\mathcal{B}(M\oplus N)(1))=\operatorname{ad}(\mathcal{B}(N))(M)=L$. Hence $K$ is generated its degree one component $K(1)$.
@@
$\Delta_{\mathcal{B}(M\oplus N)}(u)=(\operatorname{id}\#\varepsilon\otimes\operatorname{id})\Delta(u)=1\otimes u+u\otimes 1,$
@@ 4.4 From the Nichols algebra back to the space of coinvariants
Using a method similar to that in Lemma 4.3, one verifies that $W_{i}^{\prime}$ is an $\mathbb{Z}$-graded $\mathcal{A}(N)$-subcomodule. Moreover, $\mathcal{A}(N)\rhd(\mathcal{B}(N)\rhd M_{i}^{\prime})\subseteq\mathcal{A}(N)\rhd M_{i}^{\prime}=\mathcal{B}(N)\rhd M_{i}^{\prime}$. Therefore $W_{i}^{\prime}$ is a $\mathbb{Z}$-graded subobject of $W_{i}$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. The object $W_{i}^{\prime}$ is concentrated in degrees $\geq m_{i}$, and its degree $m_{i}$’s component is exactly $M_{i}^{\prime}$. Since $W_{i}$ is irreducible, we must have $W_{i}^{\prime}=W_{i}$. Therefore, $M_{i}^{\prime}=W_{i}^{\prime}\cap M=W_{i}\cap M=M_{i}$, which establishes the irreducibility of $M_{i}$. Moreover, $M_{i}$ is indeed the homogeneous component of $W_{i}$ of minimal degree. Finally, for each $i\in I$ and $n\in\mathbb{N}_{0}$,
@@
is a $\mathbb{N}_{0}$-graded Hopf algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}. Its degree 0’s part is $\mathbbm{k}$, and its degree $1$’s part is $M\oplus N$.

### v2

@@ 2.4 Nichols algebras in a braided monoidal category
Let $V$ be a nonzero object in $\mathcal{C}$; by $T(V)$ we denote the tensor algebra in $\mathcal{C}$ generated freely by $V$. Here we denote that
@@
[Meir, 2022, Definition 2.4] For a given object $V\in\mathcal{C}$ the Nichols algebra $\mathcal{B}(V)$ is the unique Hopf algebra in $\mathcal{C}$ that satisfies the following conditions:
@@
Recall that the Nichols algebras in $\mathcal{C}$ are of the form $T(X)/I$ where $X\in\mathcal{C}$ and $I$ is the unique maximal graded Hopf ideal in $T(V)$ generated by homogeneous elements of degree greater than or equal to $2$. According to $F(T(V))\cong T(F(V))$, we have
@@
which is finite-dimensional as well. Here $F(I)$ is a homogeneous Hopf ideal of degree greater than or equal to $2$ of $T(F(V))$. Note that the Nichols algebra generated by $F(V)$ must be of the form $T(F(V))/J$, where $J$ is the unique maximal homogeneous graded Hopf ideal of $T(F(V))\in\mathcal{D}$ with degree greater than or equal to $2$. Hence $F(I)\subset J$ and $T(F(V))/J\subset T(F(V))/F(I)$. On the other hand, $F^{-1}:\mathcal{D}\longrightarrow\mathcal{C}$ is the inverse of $F$, which is an exact monoidal functor. Hence
@@ 3 Monoidal equivalence between Yetter-Drinfeld module categories
We build upon the theory of Hopf pairings and partial dualization in braided monoidal categories Barvels et al. (2015). We apply this theory to the setting of coquasi-Hopf algebras. Our main result, Theorem 3.8, establishes a key braided monoidal equivalence $\Omega$ between categories of Yetter-Drinfeld modules, which is important to our reflection theory.
@@ 3.1 Hopf pairings in braided monoidal categories
In this subsection, $\mathcal{C}$ always represents a braided monoidal category with associative isomorphism $a$ and braiding isomorphism $c$ unless stated otherwise. We begin by introducing the notion of Hopf pairings.
@@
A Hopf pairing $\omega:A\otimes B\to 1$ is called non-degenerate if there is a morphism $\omega^{\prime}:1\rightarrow B\otimes A$, such that
@@
[Barvels et al., 2015, Theorem 3.20] Let $\omega:A\otimes B\to\mathbf{1}$ be a non-degenerate Hopf pairing. The categories $^{A}_{A}\mathcal{YD}(\mathcal{C})$} and $^{B}_{B}\mathcal{YD}(\mathcal{C})$} are braided monoidal equivalent via the monoidal functors $\Omega^{\omega}$ as follows.
@@
    is isomorphic to the identity functor by [Barvels et al., 2015, Proposition 3.22].
@@
    For completeness, we recall the construction of the functor $\Omega^{\omega}$, which is investigated in [Barvels et al., 2015, Theorem 3.16, Corollary 3.19]. Let $\omega:A\otimes B\to\mathbf{1}$ be a non-degenerate Hopf pairing. The functor $\Omega^{\omega}$ is given by
@@ 3.2 Partial dualization of a Hopf algebra in braided monoidal categories
We now turn our attention to the construction of partial dualization. This construction, originally developed in [Barvels et al., 2015, Theorem 4.4], provides a powerful tool for studying Hopf algebras in braided monoidal categories.
@@
We establish this result in several steps. First, the object $\mathcal{B}(N)^{*\operatorname{op,cop}}$ with the following operations is a Hopf algebra in $^{H}_{H}\mathcal{YD}$} by [Andruskiewitsch and Graña, 1999, Lemma 2.2.3]
@@
in $^{H}_{H}\mathcal{YD}$}. The object $\mathcal{B}(N^{*})$ is the graded dual of $\mathcal{B}(N)$ by [Meir, 2022, Lemma 2.6]. Therefore $\operatorname{dim}(\mathcal{B}(N^{*}))=\operatorname{dim}(\mathcal{B}(N)^{*})$ since $\operatorname{dim}(\mathcal{B}(N))$ is finite-dimensional. As a result
@@
Let $H$ be a coquasi-Hopf algebra with bijective antipode, and $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. Suppose $N=\bigoplus_{i\in I}N_{i}\in{{}^{H}_{H}\mathcal{YD}}$, and the Nichols algebras $\mathcal{B}(N_{i})$ are all finite-dimensional. For each $i\in I$, we have a partial dualization datum $\mathcal{A}_{i}=(\mathcal{B}(N)\overset{\pi_{i}}{\rightarrow}\mathcal{B}(N_{i}),\mathcal{B}(N_{i}^{*}),\operatorname{ev}_{\mathcal{B}(N_{i})}^{-})$. We set $K_{i}:=\mathcal{B}(N)^{\operatorname{co}\mathcal{B}(N_{i})}\in{{}^{\mathcal{B}(N_{i})}_{\mathcal{B}(N_{i})}\mathcal{YD}(\mathcal{C})}$. Then we have the following equivalences of braided monoidal categories
@@ 3.3 The functor $\Omega^{\omega}$ under $\mathbb{Z}$-gradings
(1) The module and comodule maps of $K(1)$ are $\mathbb{Z}$-graded and hence the infinitesimal braiding $c\in\mathrm{Aut}(K(1)\otimes K(1))$, being determined by the module and comodule maps, is $\mathbb{Z}$-graded. Moreover, the structure maps of $K(n)$ for $n\in\mathbb{N}$ are determined by $c$ and $K(1)$, since $K(1)$ generates $K$. Therefore $K$ is a $\mathbb{Z}$-graded objects in ${{}^{H}_{H}\mathcal{YD}}(\mathbb{Z}\operatorname{-Gr}\mathcal{M}_{\mathbbm{k}})$.
@@
Let $H$ be a coquasi-Hopf algebra with bijective antipode, and $\mathcal{C}={{}^{H}_{H}\mathcal{YD}}$. Suppose $R$ is an $\mathbb{N}_{0}$-graded Hopf algebra in $^{H}_{H}\mathcal{YD}$}. Let $X$ be an $R$-module in $\mathcal{C}$ and $Y$ be an $R$-comodule in $\mathcal{C}$. For $n\geq 0$, we define:
@@
With the above notations, the functor $\Omega$ relates the operator $\mathcal{F}_{n}$ to $\mathcal{F}^{n}$.
@@
Define an auxiliary map $\rho^{\prime}:=(\operatorname{ev}\otimes\operatorname{id}_{V})\circ a^{-1}_{\mathcal{B}(N^{*}),\mathcal{B}(N),V}\circ(\operatorname{id}_{\mathcal{B}(N^{*})}\otimes\delta_{\mathcal{B}(N)})$, which is derived from $\rho_{\mathcal{B}(N^{*})}$. By [Heckenberger and Schneider, [2020] ©2020, Corollary 3.3.4, Lemma 12.2.11], $(V,\rho^{\prime})$ is still a $\mathcal{B}(N^{*})$-module in $\mathcal{C}$ and
@@
is $\bigoplus_{m\neq n}\mathcal{B}(N^{*})(m)\otimes V$. This means, if $v\in\mathcal{F}^{n}V$, then for each $i>n$, $\rho^{\prime}(\mathcal{B}(N^{*})(i)\otimes v)=0$ by (3.5). On the other hand, if $v\in\mathcal{F}_{n}(V,\rho^{\prime})$, this will imply $\delta(v)\in\bigoplus_{i=0}^{n}\mathcal{B}(N)(i)\otimes V$, which completes the proof of (1): $\mathcal{F}_{n}\Omega(V)=\mathcal{F}_{n}(V,\rho^{\prime})=\mathcal{F}^{n}V$.
@@
We denote $\delta^{\prime}=(\operatorname{id}_{\mathcal{B}(N^{*})}\otimes\rho_{\mathcal{B}(N)})\circ a_{{\mathcal{B}(N^{*})},{\mathcal{B}(N)},V}\circ((\operatorname{ev}^{-})^{\prime}\otimes\operatorname{id}_{V})$, which origins from $\delta_{\mathcal{B}(N)}$. Again by [Heckenberger and Schneider, [2020] ©2020, Corollary 3.3.6, Lemma 12.2.11]
@@ 4 Projections of Nichols algebras
Let $(\mathbbm{k}G,\Phi)$ be a pointed coquasi-Hopf algebra, where $G$ is a finite group, $\Phi$ is a $3$-cocycle on $G$. In this section, we study projection of Nichols algebras in $^{G}_{G}\mathcal{YD}^{\Phi}$}, which will play an important role when considering reflection of simple Yetter-Drinfeld modules. The proofs of some lemmas and theorems in this section, while are conceptually parallel to those in the Hopf algebra setting Heckenberger and Schneider (2013a), take a different path due to the non-strict monoidal structure.
@@ 4.1 The structure of the space of coinvariant $K$
To study the projection properties of Nichols algebras, we must first examine their structure in the category $^{G}_{G}\mathcal{YD}^{\Phi}$}.
@@ 4.2 The semisimplicity of the object $L$
[Heckenberger and Schneider, [2020] ©2020, Lemma 1.3.13, Proposition 1.3.14]
@@
[Heckenberger and Schneider, [2020] ©2020,Proposition 13.2.3] Let $C$ be an $\mathbb{N}_{0}$-graded coalgebra, $X$ an $\mathbb{N}_{0}$-graded left $C$-comodule, and $Y$ a $C$-subcomodule of $X$. Let $k>0$ be an integer. Assume that the map $\delta_{n-k,k}^{X}:X(n)\to C(n-k)\otimes X(k)$ is injective for all $n\geq k$, and $Y$ is not contained in $\bigoplus_{i=0}^{k-1}X(i)$. Then $Y\cap\bigoplus_{i=0}^{k}X(i)\neq 0$.
@@
Now let $X\in L$, then $\delta(X)\in\mathcal{B}(N)\otimes L$ by Lemma 4.4, which implies that L inherits an $\mathbb{N}_{0}$-graded $\mathcal{B}(N)$-comodule structure via the restriction of $\delta:L\rightarrow\mathcal{B}(N)\otimes L$.
@@ 4.3 Realizing $K$ as a Nichols algebra
We endow a non-standard grading on $\mathcal{B}(M\oplus N)$ via setting $\operatorname{deg}(M)=1$ and $\operatorname{deg}(N)=0$. Under this grading, $\mathcal{B}(M\oplus N)$ remains a $\mathbb{N}_{0}$-graded Hopf algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}. Furthermore, $\mathcal{B}(M\oplus N)\#\mathbbm{k}G$ becomes a $\mathbb{N}_{0}$-graded coquasi-Hopf algebra with $\operatorname{deg}(\mathbbm{k}G)=0$. It follows direct that $K$ inherits an $\mathbb{N}_{0}$-graded Hopf algebra in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$} with $K(n)=K\cap(\mathcal{B}(M\oplus N)\#\mathbbm{k}G)(n)$. In particular, we have $K(1)=K\cap(\mathcal{B}(M\oplus N)\#\mathbbm{k}G)(1)=K\cap(\mathcal{B}(M\oplus N)(1))=\operatorname{ad}(\mathcal{B}(N))(M)=L$. Hence $K$ is generated by its degree one component $K(1)$.
@@
$\Delta_{\mathcal{B}(M\oplus N)}(u)=(\operatorname{id}\#\varepsilon\otimes\operatorname{id})\Delta_{\mathcal{A}(M\oplus N)}(u)=1\otimes u+u\otimes 1,$
@@ 4.4 From the Nichols algebra back to the space of coinvariants
Using a method similar to that in Lemma 4.3, one verifies that $W_{i}^{\prime}$ is a $\mathbb{Z}$-graded $\mathcal{A}(N)$-subcomodule. Moreover, $\mathcal{A}(N)\rhd(\mathcal{B}(N)\rhd M_{i}^{\prime})\subseteq\mathcal{A}(N)\rhd M_{i}^{\prime}=\mathcal{B}(N)\rhd M_{i}^{\prime}$. Therefore $W_{i}^{\prime}$ is a $\mathbb{Z}$-graded subobject of $W_{i}$ in $^{\mathcal{A}(N)}_{\mathcal{A}(N)}\mathcal{YD}$}. The object $W_{i}^{\prime}$ is concentrated in degrees $\geq m_{i}$, and its degree $m_{i}$’s component is exactly $M_{i}^{\prime}$. Since $W_{i}$ is irreducible, we must have $W_{i}^{\prime}=W_{i}$. Therefore, $M_{i}^{\prime}=W_{i}^{\prime}\cap M=W_{i}\cap M=M_{i}$, which establishes the irreducibility of $M_{i}$. Moreover, $M_{i}$ is indeed the homogeneous component of $W_{i}$ of minimal degree. Finally, for each $i\in I$ and $n\in\mathbb{N}_{0}$,
@@
is a $\mathbb{N}_{0}$-graded Hopf algebra in $^{G}_{G}\mathcal{YD}^{\Phi}$}. Its degree 0 component is $\mathbbm{k}$, and its degree $1$ component is $M\oplus N$.